# Interactive, and Editable Simulation for Autonomous Driving

### v1

@@ V Conclusion
This paper presented a novel, unified simulation framework designed to meet the multifaceted demands of autonomous driving research by combining dynamical fidelity, photorealistic rendering, scenario orchestration, and real-time performance. We demonstrated how the proposed framework effectively bridges the gap between real2sim fidelity and serviceability in digital twin simulations by integrating physics-based modeling with data-driven reconstruction and enabling natural language interaction through an LLM interface. Our experiments indicated high levels of structural similarity (97%) in reconstructed scenes/assets, strong generalizability (95%) and repeatability (85%) in scenario reconfiguration, and real-time performance ($>$60 FPS) on a consumer-grade PC. These findings suggest that the proposed framework can serve as a scalable tool for advancing the development, optimization, and validation of autonomous driving algorithms under realistic and diverse conditions.
Future research could focus on defining data requirements (quality and quantity) to further streamline the scene/asset reconstruction pipeline. Additionally, expanding the framework’s library of reconstructed scenes and assets would be a natural step towards enriching the creativity and variability of orchestrated simulations. Finally, there is a scope for enhancing the scenario reconfiguration pipeline by redefining the agentic AI topology, introducing additional context using techniques such as retrieval-augmented generation (RAG), or specializing it through fine-tuning.

### v2

@@ V Conclusion
This paper presented a novel, unified simulation framework designed to meet the multifaceted demands of autonomous driving research by combining dynamical fidelity, photorealistic rendering, context-aware scenario orchestration, and real-time performance. We demonstrated how the proposed framework effectively bridges the gap between real2sim fidelity and serviceability in digital twin simulations by integrating physics-based modeling with data-driven reconstruction, and enabling natural language interaction through an LLM interface. Our experiments indicated high levels of structural similarity ($\sim$97%) in reconstructed scenes/assets, strong generalizability ($\sim$85%) and repeatability ($\sim$95%) in scenario reconfiguration, and real-time performance (>>60 FPS) on a consumer-grade PC. Furthermore, the optional VLM-based image enhancement module consistently improved the output image quality ($\sim$80%) by seamlessly blending the scene composition. These findings suggest that the proposed framework can serve as a scalable tool for advancing the development, optimization, and validation of autonomous driving systems under realistic and diverse conditions.
Looking ahead, we wish to address some of the limitations of the existing pipeline: (a) 3DGS does not fully support raytracing, which limits lighting effects such as shadows, reflections, etc.; (b) the demonstrated LLMs tend to get stuck at certain prompts without retaining the keywords and require specific prompt engineering to circumvent this issue, which cannot be guaranteed to work every time; and (c) The VLM-based enhancement pipeline is susceptible to generating varied outputs that may, occasionally, not completely adhere to the specified constraints.